OpenStudy (anonymous):

Find all the values of t between 0 and 2pie for which sec(2t-1) = 1.25

Parth (parthkohli):

Now take \(\cos^{-1}\) (also known as \(\arccos\))

Parth (parthkohli):

Am I right?

OpenStudy (anonymous):

something looks wrong :) \[ \frac{1}{cos(2t-1)} = 1.25\] your use of notations is wrong :)

Parth (parthkohli):

Oopsie

OpenStudy (anonymous):

you cant say \(sec^{-1}\)

OpenStudy (anonymous):

or \(cos^-1\)

Parth (parthkohli):

I am still a learner, sorry.

Parth (parthkohli):

\(1 = 1.25 \cos(2t - 1)\) \[ 0.8 = \cos(2t - 1) \] Now is that correct?

OpenStudy (anonymous):

you should know what \(cos^{-1}x\) is you are NOT allowed to say \(cos^{-1}\)

Parth (parthkohli):

Hmm. I just meant 'taking arccos'

OpenStudy (anonymous):

inverse functions are different dont be confused :) You are indeed right now.

OpenStudy (anonymous):

in mathematics terminology you cant say it like that

Parth (parthkohli):

Thank you!

OpenStudy (anonymous):

\(secx\) is NOT \(cos^{-1}\)

Parth (parthkohli):

\( \arccos(0.8) = \arccos{(\cos(2t - 1))}\)

OpenStudy (anonymous):

no no

Parth (parthkohli):

Oh

Parth (parthkohli):

I need to open my own question before I spoil this one up.

OpenStudy (anonymous):

\[(2t-1) =arccos(0.8)\]

OpenStudy (anonymous):

@Omniscience and what is \[\cos^{-1}\]??

OpenStudy (anonymous):

@ParthKohli: used it wrongly :) that is not right. \[secx = \frac{1}{cosx}\] you CANNOT say its \(cos^{-1}\) it just doesnot exist :)

Parth (parthkohli):

@Omniscience By \(\cos^{-1}\), I meant 'arccos', not \(1 \over \cos\)

OpenStudy (anonymous):

well for inverse trig symbols like \(cos^{-1}\) are not allowed to stand alone because they do not make sense :) if you are referring to \(arccos\) then its fine guess but can be confusing in mathematics; better not to use it often :)

Parth (parthkohli):

Haha :)

OpenStudy (anonymous):

Now ! I agree ! @Omniscience Thank you !

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so there is a prob with the question?